Problem 9 Intro You took out a loan to buy a new car. The monthly interest rate on the loan is 2.5%. You have to pay $280 every month for 60 months. Part 1 B Attempt 1/10 for 8 pts. What is the present value of the cash flows if its an ordinary annuity? No decimals Submit Part 2 What is the future value of the cash flows if its an ordinary annuity? 8 - Attempt 1 /10 for 8 pts. No decimals Submit
Part 3 What is the present value of the cash flows if its an annuity due? ~ Attempt 1 /10 for 10 pts. No decimals Submit Part 4 What is the future value of the cash flows if its an annuity due? ~ Attempt 1 /10 for 8 pts. No decimals Submit
Answer #1

Answer Part 1:

Monthly Interest rate = 2.5%

Monthly payment = $280

Period = 60 months

Ordinary annuity

To get present value we will use PV function of excel:

= PV (rate, nper, pmt, fv, type)

= PV (2.5%, 60, -280, 0, 0)

= $8654.424

$8,654

Answer Part 2:

Future value of ordinary annuity

We will use FV of excel:

= FV (rate, nper, pmt, pv, type)

= FV (2.5%, 60, -280, 0, 0)

= $38,077.645

$38,078

Answer Part 3:

Present value of annuity due

= PV (2.5%, 60, -280, 0, 1)

= $8870.784

$8,871

Answer Part 4:

Future value of annuity due

= FV (2.5%, 60, -280, 0, 1)

= $39029.586

$39,030